OpenStudy (agentnao):

Begin with one cent on day 1, four cents on day 2, 16 cents on day three, and so on until the end of 18 days. How much money will there be total?

OpenStudy (akshay_budhkar):

imran?

OpenStudy (agentnao):

Shouldn't that be 18 on top and 1 on the bottom?

OpenStudy (akshay_budhkar):

16 on 3 imran

OpenStudy (agentnao):

And why is it n^2? I thought it would be 4n?

OpenStudy (akshay_budhkar):

no no neither i believe

OpenStudy (anonymous):

not clear is it? it could be \[4^k\] as well

OpenStudy (anonymous):

\[1,4,16,64,...\]

OpenStudy (agentnao):

Would choosing this option be more money than 100,000,000?

OpenStudy (anonymous):

not \[2^k\] for sure

OpenStudy (amistre64):

n: 1 2 3 4 A: 1 4 16 ? 4^(n-1) right?

OpenStudy (anonymous):

i think it is \[\sum_{k=0}^{18}4^k\]

OpenStudy (akshay_budhkar):

yes i think satellite is right

OpenStudy (anonymous):

you can start at zero and write what i wrote or start at one and write amsitre's

OpenStudy (across):

\[\sum_{i = 0}^{17}4^i\]

OpenStudy (amistre64):

have we written out a sequence for the partial sums yet? or am i readin ghtis wrong?

OpenStudy (anonymous):

anyway it is a huge number because \[4^{17}=17,179,869,184\]

OpenStudy (anonymous):

the total is \[\frac{4^{18}-1}{4-1}\] whatever that is

OpenStudy (agentnao):

Is the total 229064922 by anychance?

OpenStudy (anonymous):

22906492245

OpenStudy (anonymous):

isnt it just geometric series

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

there's a formula for this just plug it in

OpenStudy (agentnao):

It is "just geometric series", but it's not working out as planned.

OpenStudy (agentnao):

I tried that already. I wouldn't be asking if it were that simple.

OpenStudy (anonymous):

do you perhapsknow the final answer?

OpenStudy (agentnao):

Wait, satellite! Because this is in pennies, wouldn't we have to divide by 100 to convert to dollars?

OpenStudy (agentnao):

No hahd, not yet :(

OpenStudy (anonymous):

yes you would

OpenStudy (agentnao):

Which would then make it LESS than 100,000,000, right?

OpenStudy (anonymous):

the formula for summing a geometric series is the one i wrote. it is \[\frac{a(r^{n+1}-1)}{r-1}\] in this case r is 4 and a is 1 i believe.

OpenStudy (agentnao):

A lottery winner must decide between two methods of payment. Choice one is to receive a lump sum of $100,000,000. Choice two is to begin with one cent on day 1, four cents on day 2, 16 cents on day three, and so on until the end of 18 days. Which choice should the winner pick. <-- that is the whole question.

OpenStudy (agentnao):

I thought a would be .01 since it is a penny. and why is it n + 1? I thought it was n - 1

OpenStudy (across):

Isn't the final answer $229,064,922.45?

OpenStudy (anonymous):

end of 18 days maybe means you get that too, so it could be \[\frac{4^{19}-1}{3}\] take the series, you will make more for sure

OpenStudy (agentnao):

across, that's what I thought, but I think you'd have to divide that by 100 to convert it to dollars because what you have is in pennies.

OpenStudy (across):

I already divided it by 100. ^^

OpenStudy (anonymous):

give me a second ill writer a program

OpenStudy (amistre64):

another form of it is: i think \[\frac{1-r^n}{1-r}\]

OpenStudy (agentnao):

But that can't be. That's the same answer I got without dividing by 100.

OpenStudy (agentnao):

http://www.twiddla.com/596084 Let's work it out in here.

OpenStudy (amistre64):

22,906,492,245 pennies is what I get

OpenStudy (agentnao):

Really? How did you get that? Show me in the whiteboard room.

OpenStudy (amistre64):

lets say the amount = S \[S = 4^0+4^1+4^2+...+4^{n-1}\]now mulitply it all by 4 to get \(4S =\) \( 4^1+4^2+...+4^{n-1}+4^n\) subtract them: S =\( 4^0+4^1+4^2+...+4^{n-1}\) -4S = \( -4^1-4^2-...-4^{n-1}-4^n\) ------------------------------------ (1-4)S =\(4^0\) \(-4^n\) now divide off the (1-4) to get S = \(\cfrac{1-4^n}{1-4}\)

OpenStudy (agentnao):

And then just plug in 18 for n and I should get my answer?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

thats your answer in pennies :)

OpenStudy (agentnao):

Thank you very much amistre!

OpenStudy (amistre64):

1 penny = .01 dollars youre welcome :)